Hoodmemes:

If Jacob is rollerskating for 1.5 miles and it takes him 17 minutes, what was his average speed?

jhonyy9:

use the rule of 3 simple and write 1,5 miles in 17 min. x miles in 1 min. cross multiplie and calcule the x value x = 1*1,5/17 = ? mil

Hoodmemes:

"1*1,5/17" what does the 5/17 mean?

jhonyy9:

there is 1,5 /17 = ?

jhonyy9:

in this way will get how many miles will take in 1 min.

Hoodmemes:

I'm not sure what / meaning is it divide or multiply?

jhonyy9:

divide

Hoodmemes:

Well I use calculator and is said 0.08823529411

jhonyy9:

what mean this ?

Hoodmemes:

It said 1.5 divided by is 17 was 0.08823529411

jhonyy9:

that he take 0,088 miles in 1 min.

Hoodmemes:

Oh I needed to round